Calculate the Median and Mode from the following data:
Annual Sales(Rs.000) Frequency
Less than 10 4
Less than 20 20
Less than 30 35
Less than 40 55
Less than 50 62
Less than 60 67
1
Expert's answer
2020-12-10T13:02:35-0500

Solution

Median:

"Median= L+( {{N \over 2}-cf \over f})*i"

"{N \over 2}= {67 \over 2}= 33.5 \space \therefore" the median class is 20-29


"Median = 19.5+ ({{67 \over 2} - 20 \over 15})*10"


=28.5

=Rs 28500


Mode:

"Mode = L+({f_1-f_0 \over 2f_1-f_0-f_1})*i"

30-39 is the modal class since it has the highest frequency


"Mode= 19.5+({20-15 \over 2(20)-15-7})*10"

"=29.5+2.77778"

=32.27778

=Rs. 32277.8